1. On January 1, 2020, Breton Company issued its 8% bonds in the face amount of $3,000,000, which mature on January 1, 2030. The bonds were issued for $3,441,591 to yield 6%. Geller uses the effective-interest method of amortizing bond premium. Interest is payable annually on December 31. Interest Expense for 2023 is:
Answer
$_______________
2. On May 1, 2020, Judice Company issued 400 $1,000 bonds at 104. Each bond was issued with two detachable stock warrants. Shortly after issuance, the bonds were selling at 106, and the fair value of the warrants was $40 each.
Prepare the entry to record the issuance of the bonds and warrants.
Answer #1

Solution:

(1) Computation of Interest Expenses for 2023:

Beginning of Period Net Premium End of Period Interest Amount of Book Value Amortizati Period Net Ending Expenses Payment of

Interest expense for the Year 2023 will be $ 200,095.56 approximately.

(2) Journal Entry to record issuance of bonds with warrants:

In this question, we will use the proportional method to appropriate proceeds into bond and warrant.

Computation of Fair Value Proportions:

Total Fair Value of bond     = 106 + (2x40) = 186

Proportion for Bond part     = 106/186

Proportion for warrant part = 80/186

Appropriation of proceeds:

Total Proceeds         =          400 x $1000 x 104   =          $4,16,00,000

Bond part                   =          $4,16,00,000 x 106/186      =          $2,37,07,526.89

Warrant Part              =          $4,16,00,000 x 80/186         =          $1,78,92,473.11

Journal Entry:

            Bank A/c Dr.                          $4,16,00,000

                        To Bonds payable A/c                                             $2,37,07,526.89

                        To Paid in capital – Stock Warrants A/c               $1,78,92,473.11

            (Being amount received on issuance of bond with warrants)